# LLaVA-MoD: Making LLaVA Tiny via MoE-Knowledge Distillation

### v1

@@ Abstract
We introduce LLaVA-MoD, a novel framework designed to enable the efficient training of small-scale Multimodal Language Models (s-MLLM) by distilling knowledge from large-scale MLLM (l-MLLM). Our approach tackles two fundamental challenges in MLLM distillation. First, we optimize the network structure of s-MLLM by integrating a sparse Mixture of Experts (MoE) architecture into the language model, striking a balance between computational efficiency and model expressiveness. Second, we propose a progressive knowledge transfer strategy to ensure comprehensive knowledge migration. This strategy begins with mimic distillation, where we minimize the Kullback-Leibler (KL) divergence between output distributions to enable the student model to emulate the teacher network’s understanding. Following this, we introduce preference distillation via Direct Preference Optimization (DPO), where the key lies in treating l-MLLM as the reference model. During this phase, the s-MLLM’s ability to discriminate between superior and inferior examples is significantly enhanced beyond l-MLLM, leading to a better student that surpasses its teacher, particularly in hallucination benchmarks. Extensive experiments demonstrate that LLaVA-MoD outperforms existing models across various multimodal benchmarks while maintaining a minimal number of activated parameters and low computational costs. Remarkably, LLaVA-MoD, with only 2B activated parameters, surpasses Qwen-VL-Chat-7B by an average of 8.8% across benchmarks, using merely $0.3\%$ of the training data and 23% trainable parameters. These results underscore LLaVA-MoD’s ability to effectively distill comprehensive knowledge from its teacher model, paving the way for the development of more efficient MLLMs. The code will be available on https://github.com/shufangxun/LLaVA-MoD
Figure 1: Comparisons of training cost and performance. LLaVA-MoD achieves competitive performance with advanced MLLMs using significantly lower training costs while outperforming current small-scale MLLMs by a large margin.
@@ 1 Introduction
Leveraging the advanced instruction-following and reasoning capabilities of Large Language Models (LLMs) (Achiam et al., 2023; Bai et al., 2023a; Jiang et al., 2024; Team et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024), Multimodal Large Language Models (MLLMs) (Bai et al., 2023b; Liu et al., 2024; Lin et al., 2024b; Li et al., 2023b; Chen et al., 2023b; Shu et al., 2023; Lu et al., 2024), typically integrate a visual encoder (Caron et al., 2021; Radford et al., 2021; Liu et al., 2022) alongside an LLM to achieve promising results across various vision and cross-modal tasks, such as image captioning and visual question-answering. MLLMs are characterized by their large model size and extensive training data, which contribute to their superior performance but also necessitate significant computation resources. For instance, the largest version of LLaVA-NeXT (Li et al., 2024a) employs Qwen-1.5-110B (Yang et al., 2024) as the language model and requires 18 hours of training with 128 H800 GPUs. Furthermore, MLLMs face deployment challenges due to the high memory and computational demands. For instance, models with substantial parameters often require advanced hardware and exhibit slow inference speeds, which significantly hinder their real-world applications such as on mobile devices. Thus, developing MLLMs that balance performance and efficiency has become a crucial research focus.
@@
Previous works on small-scale Multimodal Large Language Models (s-MLLM) (Zhou et al., 2024a; Yuan et al., 2023; Shao et al., 2024; He et al., 2024; Chu et al., 2023; 2024; Yao et al., 2024) have mainly concentrated on devising data collection (Schuhmann et al., 2021; 2022; Zhu et al., 2024; Awadalla et al., 2024; Gadre et al., 2024; Chen et al., 2023a) and filtering protocols (Xu et al., 2023; Fan et al., 2024; Zhang et al., 2023; Li et al., 2024b; Nguyen et al., 2024) to ensure high-quality training data. While such high-quality data aid s-MLLM in narrowing the performance gap with their larger counterparts, their effectiveness is still limited by model capacity and the availability of high-quality data. With the advent of open-source MLLMs, an intuitive strategy inspired by conventional deep learning tasks involves utilizing large-scale MLLMs (l-MLLM) as the teacher to improve the training of s-MLLM via Knowledge Distillation (KD) (Hinton et al., 2015). KD facilitates stable and efficient training of smaller models by aligning the output distributions of the student model with those of the teacher model. However, the application of KD as a model reduction technique has yet to be extensively explored in MLLMs. To develop an effective framework for MLLM distillation, we consider two primary challenges. The first challenge lies in designing a lightweight architecture for student MLLMs that retains strong learning and expressive capabilities, enabling it to absorb the complex knowledge embedded within teacher MLLMs effectively. The second challenge is to efficiently and comprehensively transfer this knowledge from the teacher MLLMs to the student MLLMs.
@@
To address these two challenges, we present LLaVA-MoD [^1] MoD denotes Mixture-of-Expert Knowledge Distillation for MLLM, an effective framework for training s-MLLM by mimicking the behavior of the l-MLLM through Mixture-of-Expert (MoE) Knowledge Distillation. For the first challenge—designing a lightweight s-MLLM structure—a straightforward approach might involve reducing the scale of the base language model in l-MLLM to create a smaller network. However, this direct reduction in network parameters can significantly compromise the model’s expressive capability, making it less effective in modeling complex multimodal understanding tasks. Drawing inspiration from the recent success of sparse MoE (Lin et al., 2024a; Dai et al., 2024; Jiang et al., 2024; Shen et al., 2023) in sequential modeling, we incorporate a sparse MoE structure into the dense s-MLLM. This approach seeks to balance scale reduction while preserving the model’s ability to capture and represent intricate multimodal information during distillation. Specifically, we enhance the s-MLLM by equipping it with several feedforward networks (FFNs) and linear gates within the LLM. Each FFN acts as an expert, capturing fine-grained knowledge from the l-MLLM, while the gates select the top-$k$ experts to facilitate optimal knowledge transfer pathways.
@@
To address the second challenge, we propose a progressive distillation strategy for effective knowledge transfer. The process begins by aligning the vision encoder with the LLM using a learnable adapter, thereby initializing a dense student MLLM. Following this, we employ two consecutive distillation stages, where the student model evolves from mimicking and approximating the teacher MLLM to ultimately surpassing it: Mimic Distillation. This stage is divided into two steps, i.e., dense-to-dense and dense-to-sparse distillation. In the dense-to-dense distillation stage, we employ standard knowledge distillation loss to align the output logits distribution between the initialized dense s-MLLM and the teacher model, using general captioning and conversation datasets. Next, we transition to the dense-to-sparse distillation where we transform the dense student MLLM into a sparse one by integrating MoE, and then distill knowledge from the teacher MLLM into this sparse student model using a broad range of tasks and datasets. Preference Distillation. In this stage, the teacher model provides knowledge regarding what constitutes “good” and “bad” samples, establishing a foundational reference for the student model. The student MLLM leverages this knowledge to adjust its probability distribution, ensuring that good samples have a higher probability than those from the teacher model, while bad samples are assigned a lower probability. This process enhances the student model’s ability to mitigate hallucinations by improving its judgment capabilities beyond those of the teacher model.
@@
As illustrated in Figure 1, LLaVA-MoD exhibits impressive performance across various multimodal benchmarks while maintaining minimal activated parameters and low computational resources. For example, LLaVA-MoD-2B exceeds Qwen-VL-Chat-7B by an average of $8.8\%$ on these benchmarks, utilizing just $0.3\%$ of the training data and $23\%$ trainable parameters. Furthermore, it matches the performance of RLHF-based methods with 7B and 13B parameters on several hallucination benchmarks. Specifically, LLaVA-MoD-2B surpasses RLHF-V (Yu et al., 2024a) by $8.2\%$ in response-level hallucination rate and by $21.3\%$ in mention-level hallucination rate on the Object HalBench. The impressive results demonstrate the effectiveness of our proposed MoD framework in transferring knowledge from l-MLLM to s-MLLM.
@@ 2 Related Work
Multimodal Large Language Models. The emergence of large language models (LLMs) has greatly advanced the field of natural language processing. Connecting visual information to LLM to enhance their understanding of multimodal input is crucial for promoting the unified comprehension of vision and language. CLIP (Radford et al., 2021) was the first to align visual and textual information into a unified embedding space through contrastive learning objectives. BLIP-2 (Li et al., 2023b) utilizes various pre-training tasks to train separate visual and text streams, and adds additional intermediary structures to adapt visual features to LLM. Flamingo (Alayrac et al., 2022) incorporates additional cross-attention modules into LLM to handle arbitrarily interleaved sequences of images and texts. Recently, models like LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) use linear projection layers to map image features to the textual space and enhance the models’ instruction-following ability through visual instruction tuning. Moreover, some works focus on a stronger vision encoder. Intern-VL (Chen et al., 2023b) leverages a larger-scale vision encoder to bridge the gap with a large-scale language model, and Mini-Gemini (Li et al., 2024c) introduces additional vision encoders for high-resolution refinement to enhance visual tokens. Qwen-VL (Bai et al., 2023b), VisionLLM (Wang et al., 2023c), and others aim to imbue models with stronger fine-grained visual understanding capabilities, such as grounding and region understanding. Unlike these approaches, our method does not emphasize enlarging the model to boost multimodal capabilities. Instead, we adopt a distillation approach to compress the capabilities of larger models into a smaller, sparse MoE architecture, enhancing computational and storage efficiency.
@@
Knowledge Distillation. Large language models possess powerful capabilities, but their enormous size and high inference costs limit their application in low-resource environments. Knowledge distillation (Hinton et al., 2015) uses a large model as teacher to transfer its advanced knowledge to a smaller student model, which plays a critical role in compressing model size and enables smaller models to self-improve. Some works adopt additional designs to make the distillation process more suitable for LLMs. MiniLLM (Gu et al., 2023) minimizes the reverse KLD (Kullback–Leibler divergence) to prevent students from overestimating low-probability regions in the teacher distribution, while GKD (Agarwal et al., 2023) introduces generalized knowledge distillation and promotes the integration of distillation with RLHF (Reinforcement Learning from Human Feedback). Additionally, Huang et al. (Huang et al., 2022) introduce context distillation, and some works (Mukherjee et al., 2023; Li et al., 2022; Ho et al., 2023) adopt CoT (Chain-of-Thought) distillation to enhance specific skills of small models. Our approach innovatively distills the knowledge of MLLM into a smaller sparse MoE (Mixture-of-Experts) architecture, significantly enhancing the multimodal processing capabilities of small models at a minimal cost.
@@
Mixture-of-Experts The mixture of experts (MoE) architecture is initially introduced by Jacobs et al. (1991) to enhance model performance across various samples through the utilization of independent experts. Recently, with the rise of transformer architectures, numerous studies have extended transformer layers to incorporate MoE. In transformer-based MoE architecture, feed-forward neural network (FFN) layers are often replaced with sparsely activated experts, employing a trainable Top-k gating strategy (Lepikhin et al., 2020; Fedus et al., 2022), thereby effectively increasing model capacity while maintaining a low computational overhead. Additionally, to further mitigate training costs, the sparse upcycling method (Komatsuzaki et al., 2022) is adopted, which initializes expert parameters using those from a well-trained dense model. Currently, MoE architectures have found widespread application not only in pre-trained language models (Jiang et al., 2024; Dai et al., 2024) but also in vision models and vision-language models (Lin et al., 2024a; Shen et al., 2023). Our approach integrates MoE with knowledge distillation techniques to provide stronger signals for sparse training, which remarkably decrease the training cost associated with sparse models.
@@ 3 Method
We introduce LLaVA-MoD, a novel framework for building efficient MLLM using mixture-of-experts (MoE) knowledge distillation. Our framework consists of two main components: (a). Architecture Design of s-MLLM: As shown in Figure 3, we design the s-MLLM with a sparse MoE framework, enhancing the ability to acquire specialized expert knowledge. (b). Distillation Mechanism: We design a progressive distillation mechanism as shown in Figure 2 to transfer knowledge from the l-MLLM to the sparse s-MLLM. This process involves two stages: mimic distillation followed by preference distillation. We will introduce LLaVA-MoD in detail from these two components.
@@ 3.1 Architecture Design of Sparse s-MLLM
s-MLLM Definition.
As illustrated in Figure 3, the basic architecture of s-MLLM consists of three primary components: a vision encoder, a large language model (LLM), and a vision-language (VL) adaptor. Given a multimodal instruction conversation $(x,y)$, we define our s-MLLM to process response $y$ as follows:
@@
Figure 3: Dense to Sparse Architecture of LLaVA-MoD. The architecture of LLaVA-MoD consists of three components: VL Adaptor, Vision Encoder, and LLM. The VL Adaptor and Vision Encoder remain unchanged, while the LLM is upcycled from dense to sparse.
Mixture-of-Experts.
The principle of building our s-MLLM is downsizing the LLM while leaving the vision encoder and vision-language adaptor unmodified. To achieve this downsizing goal, we sparsify the dense s-MLLM by incorporating an MoE architecture. Specifically, Figure 3 illustrates the process, where we apply the sparse upcycling technique (Komatsuzaki et al., 2022) to replicate $N$ feedforward networks (FFNs) as the expert modules. Additionally, we introduce a linear layer as the router, which dynamically activates the appropriate experts by predicting the probability of expert assignment. Given each token $x$ in the sequence, we first compute the routing value of $N$ experts:
@@ 3.2 Progressive Distillation
Our progressive distillation consists of two distinct stages, i.e., mimic distillation (Figure 2 (a)) and preference distillation (Figure 2 (b)). In the mimic distillation stage, the student MLLM $\pi_{\text{S}}$ imitates the general and specific knowledge from the teacher MLLM $\pi_{\text{T}}$. In the preference distillation stage, the student MLLM gains the preference knowledge of the teacher MLLM to further refine its output and reduce hallucinations. We use the pre-trained l-MLLM as the teacher MLLM, which is frozen during the whole distillation process. Both teacher and student MLLMs are in the same LLM family, ensuring consistent vocabulary space, which is essential for the student MLLM to imitate the teacher MLLM accurately. For simplicity, we denote the student as $\pi_{\text{S}}$ and the teacher as $\pi_{\text{T}}$.
Initialization.
Before distillation, we first align the vision encoder with the LLM via a learnable adapter, aiming to obtain a well-initialized dense student MLLM. The $\text{LLM}_{\phi}$ and $\text{ViT}_{\chi}$ are maintained frozen since their pre-trained parameters have already captured rich visual and language knowledge. Only $\text{Proj}_{\omega}$ is optimized to bridge the gap between the vision and language domain. For the initialization, we utilize common image-caption pairs from a widely used and curated dataset, which covers a diverse range of subjects and visual entities. The training objective is to minimize the cross-entropy of the generated tokens. The objective function is defined as:
@@
Mimic Distillation.
We decompose the comprehensive knowledge of the teacher MLLM into general and expertise aspects and undertake dense-to-dense distillation and dense-to-sparse distillation respectively to transfer this knowledge to the student MLLM.
a). Dense-to-Dense Distillation. In this stage, we aim to replicate the general knowledge of the teacher MLLM. Acquiring general knowledge first is crucial, as it provides a broad foundation and common understanding across various domains, enabling the student MLLM to develop a basic framework applicable in multiple scenarios. This foundation supports a more comprehensive and flexible understanding before progressing to specialized tasks. To achieve this, we maintain $\text{ViT}_{\chi}$ frozen, and jointly optimize $\text{LLM}_{\phi}$ and $\text{Proj}_{\omega}$, with trainable parameters $\theta=\{\omega,\phi\}$. We leverage common image-caption pairs and conversation datasets. The training objective is to minimize the Kullback-Leibler divergence (KLD) between the output logits of the dense student MLLM and the teacher MLLM. The objective function is defined as:
@@
where $V$ denotes the vocabulary, while $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ and $\pi_{\text{T}}(y_{k}\mid y_{<k},x)$ denote the probability of the predicted tokens for student and teacher models, respectively.
b). Dense-to-Sparse Distillation. In this stage, our focus shifts to transferring the specialized expertise of the teacher MLLM, allowing the student MLLM to obtain advanced capabilities and achieve superior performance in diverse tasks. However, directly learning this knowledge in its dense form could lead to inefficiencies and difficulty. Therefore, we sparsify the dense student MLLM by introducing multiple experts. As detailed in Section 3.1, we replicate $N$ feedforward networks (FFNs) within $\text{LLM}_{\phi}$ and incorporate an MLP layer as a router, which jointly constitutes the experts with parameters $\phi_{e}$. This sparse architecture allows the student MLLM to selectively activate the most relevant experts for different tasks and inputs, offering significant advantages in emulating the teacher’s specialized knowledge. For training, we leverage multi-task data, updating only the experts and the adaptor. We employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. Similar to the previous stage, we also adopt the KLD as the training objective. Additionally, we include the standard next-token prediction objective, i.e., minimizing the student-generated tokens’ cross-entropy loss, to inject supervision from ground truth data which can reduce the existing bias in the teacher model. The final objective function is defined as:
@@
Preference Distillation.
In this stage, our goal is to distill the preference knowledge from the teacher MLLM to guide the student MLLM towards generating not only accurate but also reasonable responses, which is crucial in reducing hallucinations. For the training process, we make effective use of preference data, which comprises meticulously paired positive responses $y^{+}$ and negative responses $y^{-}$ for the identical prompt $x$. Our preference distillation strategy is inspired by recent advancements in Direct Preference Optimization (DPO) (Rafailov et al., 2024), which bypasses the need for training a reward model by directly training on an offline preference dataset. Our key insight is to treat the teacher MLLM as the reference model during distillation. The teacher model plays a crucial role by providing insights into what constitutes “good” and “bad”, thereby establishing a fundamental reference for the student model.
Specifically, the training objective is to optimize the student model to assign higher probabilities to positive responses than those of the teacher model while assigning lower probabilities to negative responses than those of the teacher model. This training process involves two key optimization aspects: First, the student model seeks to align with the teacher model in distinguishing positive from negative responses. Second, the student model aims to surpass the teacher model by assigning higher probabilities to positive responses and lower probabilities to negative responses. We only train the mixture of experts and the VL adaptor in the student model and employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. The object function is defined as:
@@
where $\pi_{\text{S}}(y^{+}\mid x)$ and $\pi_{\text{S}}(y^{-}\mid x)$ denote the probabilities of positive and negative responses for the student model, respectively, and $\pi_{\text{T}}(y^{+}\mid x)$ and $\pi_{\text{T}}(y^{-}\mid x)$ denote the same for the teacher model.

### v3

@@ Abstract
We introduce LLaVA-MoD, a novel framework designed to enable the efficient training of small-scale Multimodal Language Models (s-MLLM) distilling knowledge from large-scale MLLM (l-MLLM). Our approach tackles two fundamental challenges in MLLM distillation. First, we optimize the network structure of s-MLLM by integrating a sparse Mixture of Experts (MoE) architecture into the language model, striking a balance between computational efficiency and model expressiveness. Second, we propose a progressive knowledge transfer strategy for comprehensive knowledge transfer. This strategy begins with mimic distillation, where we minimize the Kullback-Leibler (KL) divergence between output distributions to enable s-MLLM to emulate l-MLLM’s understanding. Following this, we introduce preference distillation via Preference Optimization (PO), where the key lies in treating l-MLLM as the reference model. During this phase, the s-MLLM’s ability to discriminate between superior and inferior examples is significantly enhanced beyond l-MLLM, leading to a better s-MLLM that surpasses l-MLLM, particularly in hallucination benchmarks. Extensive experiments demonstrate that LLaVA-MoD surpasses existing works across various benchmarks while maintaining minimal activated parameters and low computational costs. Remarkably, LLaVA-MoD-2B surpasses Qwen-VL-Chat-7B with an average gain of 8.8%, using merely $0.3\%$ of the training data and 23% trainable parameters. The results underscore LLaVA-MoD’s ability to effectively distill comprehensive knowledge from its teacher model, paving the way for developing efficient MLLMs. The code is available at https://github.com/shufangxun/LLaVA-MoD.
@@ 1 Introduction
Multimodal Large Language Models (MLLM) (Bai et al., 2023b; Liu et al., 2024; Lin et al., 2024b; Chen et al., 2023b; Shu et al., 2023; Lu et al., 2024) have shown promising results in multimodal tasks by integrating visual encoders with Large Language Models (LLM) (Achiam et al., 2023; Bai et al., 2023a; Team et al., 2023; Touvron et al., 2023). However, the large model size and extensive training data present significant computational challenges. For example, the largest version of LLaVA-NeXT employs Qwen-1.5-110B and trains with 128 H800 GPUs for 18 hours. Additionally, the substantial number of parameters requires advanced hardware and results in slow inference speed, complicating real-world deployment, especially on mobile devices. Therefore, exploring a small-scale MLLM (s-MLLM) that balances performance and efficiency is a crucial topic.
Figure 1: Comparisons of training cost and performance. LLaVA-MoD achieves comparable performance with advanced MLLMs using significantly lower training costs while outperforming current small-scale MLLMs by a large margin.
@@
Previous works on s-MLLM (Zhou et al., 2024a; Yuan et al., 2023; Shao et al., 2024; He et al., 2024; Chu et al., 2023; Yao et al., 2024) have focused on high-quality data collection and filtering protocols. While effective, they are inherently limited by the model capacity. Knowledge Distillation (KD) from large-scale MLLM (l-MLLM) offers a promising yet unexplored strategy for enhancing s-MLLM performance. By aligning small models’ output distributions with large models, KD allows s-MLLM to leverage the rich knowledge embedded in l-MLLM. We explore a distillation framework for MLLM and consider two primary challenges. The first challenge is to design a lightweight architecture for s-MLLM while still maintaining strong learning and expressive capabilities to effectively absorb complex knowledge from l-MLLM. The second challenge is to ensure an effective transfer of complex, multi-task knowledge from l-MLLM to s-MLLM. We present LLaVA-MoD [^1] MoD denotes Mixture-of-Expert Knowledge Distillation for MLLMs to address the challenges through Mixture-of-Expert (MoE) and Knowledge Distillation.
@@
To address the first challenge, one straightforward approach is to create a small model by simply scaling down l-MLLM. However, this direct reduction in size can significantly diminish the model’s expressive capabilities, resulting in a decline in performance for handling complex multimodal tasks. Drawing inspiration from the recent success of MoE (Dai et al., 2024; Jiang et al., 2024) in language modeling, We design an MoE structure into s-MLLM to balance the scale reduction while maintaining the ability to capture complex multimodal knowledge through sparsely activated experts during distillation. Specifically, we equip s-MLLM with multiple feedforward networks (FFNs) and a linear gate within the LLM block. Each FFN serves as an expert to capture fine-grained knowledge from l-MLLM, while the gate dynamically selects top-$k$ experts to identify the optimal knowledge transfer pathway and maintain the training and inference efficiency.
@@
To address the second challenge, we propose a progressive distillation strategy. The process begins by aligning the vision encoder with the LLM using a learnable adapter, thereby initializing a dense s-MLLM. Subsequently, we employ two consecutive distillation stages, where s-MLLM evolves from mimicking and approximating l-MLLM to ultimately surpassing it: Mimic Distillation. This stage is divided into two steps, i.e., dense-to-dense (D2D) and dense-to-sparse (D2S). The rationale for this two-step process is to facilitate the transfer of complex knowledge, where the first step targets general knowledge to build a solid foundation, while the second step targets specialized knowledge to handle multi-task tasks. In D2D, we employ standard KD loss to align the output logits distribution between the initialized dense s-MLLM and l-MLLM, using general captioning and conversation datasets. Next, in D2S, s-MLLM is first transformed from dense to sparse. We then employ the standard KD loss and LM loss to distill the MoE block, using the complex multi-task datasets. Preference Distillation. In this stage, l-MLLM provides knowledge regarding what constitutes “good” and “bad” samples, establishing a foundational reference for the student model. The s-MLLM leverages this knowledge to adjust its probability distribution, ensuring that good samples have a higher probability than those from l-MLLM, while bad samples are assigned a lower probability. This process enhances the s-MLLM’s ability to mitigate hallucinations by improving its judgment capabilities beyond those of l-MLLM.
@@
LLaVA-MoD exhibits impressive performance across various multimodal benchmarks while maintaining minimal activated parameters and low computational resources. As illustrated in Figure 1, LLaVA-MoD-2B exceeds Qwen-VL-Chat-7B by an average of $8.8\%$ on these benchmarks, utilizing just $0.3\%$ of the training data and $23\%$ trainable parameters. Furthermore, it matches the performance of RLHF-based methods with 7B and 13B parameters on several hallucination benchmarks. Specifically, LLaVA-MoD-2B surpasses RLHF-V (Yu et al., 2024a) by $8.2\%$ in response-level hallucination rate and by $21.3\%$ in mention-level hallucination rate on the Object HalBench. The impressive results demonstrate the effectiveness of our distillation framework in transferring knowledge from l-MLLM to s-MLLM.
@@ 2 Related Work
Multimodal Large Language Models. LLMs have greatly advanced the field of natural language processing. Connecting visual information to LLM is crucial for promoting the unified comprehension of vision and language. BLIP-2 (Li et al., 2023b) adds additional intermediary structures to adapt visual features to LLM. Flamingo (Alayrac et al., 2022) incorporates additional cross-attention modules into LLM to handle arbitrarily interleaved multimodal sequences. Recently, models like LLaVA (Liu et al., 2024) and MiniGPT-4 (Zhu et al., 2023) have tried to enhance the models’ instruction-following ability through visual instruction tuning. In addition, some works focus on a stronger vision encoder (Chen et al., 2023b; Li et al., 2024) or more powerful fine-grained visual understanding capabilities (Bai et al., 2023b; Wang et al., 2024). Unlike these approaches, our method does not need to significantly enlarge the model. We aim to combine distillation and MoE to improve computational and storage efficiency while maintaining advanced performance.
@@
Knowledge Distillation. The enormous sizes and high inference costs of LLMs limit their application in low-resource environments. Knowledge distillation (Hinton et al., 2015) uses a large model as the teacher to transfer its advanced knowledge to a smaller student model, which plays a critical role in compressing model size and enables smaller models to self-improve. MiniLLM (Gu et al., 2023) minimizes the reverse KL divergence to prevent students from overestimating low-probability regions in the teacher distribution, while GKD (Agarwal et al., 2023) introduces generalized knowledge distillation and promotes the integration of distillation with RLHF. Additionally, some works adopt context distillation (Huang et al., 2022) and CoT (Chain-of-Thought) distillation (Mukherjee et al., 2023; Li et al., 2022; Ho et al., 2022) to enhance specific skills of small models. Our approach innovatively distills the knowledge of MLLM into a smaller sparse MoE architecture, significantly enhancing the multimodal processing capabilities of small models at a minimal cost.
@@
Mixture-of-Experts The mixture of experts (MoE) architecture, introduced by (Jacobs et al., 1991), enhances performance by leveraging independent experts to handle diverse samples. In transformer-based architecture, the feed-forward neural network (FFN) layers serve as the experts, sparsely activated through a gating strategy (Lepikhin et al., 2020; Fedus et al., 2022). This design effectively enhances model capacity while keeping computational overhead low. Additionally, the sparse upcycling method (Komatsuzaki et al., 2022) which initializes experts using those from a well-trained dense model is proposed to further mitigate training costs. MoE has shown great potential not only in language models (Jiang et al., 2024; Dai et al., 2024) but also in vision models (Riquelme et al., 2021) and vision-language models (Lin et al., 2024a; Shen et al., 2023). Our approach integrates MoE with knowledge distillation techniques to provide stronger signals for sparse training, which remarkably decreases the training cost associated with sparse models.
@@ 3 Method
We introduce LLaVA-MoD, a novel framework for building efficient s-MLLM using mixture-of-experts (MoE) and knowledge distillation. Our framework consists of two main components: (a). Architecture Design of s-MLLM: As shown in Fig. 3, we design a sparse s-MLLM with MoE, enhancing the ability to acquire specialized expert knowledge while maintaining training and inference efficiency. (b). Distillation Mechanism: We design a progressive distillation mechanism as shown in Fig. 2 to transfer knowledge from l-MLLM to sparse s-MLLM. This process involves two stages: mimic distillation followed by preference distillation.
@@ 3.1 Architecture Design of Sparse s-MLLM
s-MLLM Definition. As illustrated in Fig. 3, the basic architecture of s-MLLM consists of three primary components: a vision encoder, a large language model (LLM), and a vision-language (VL) adaptor. Given a multimodal instruction conversation $(x,y)$, we define our s-MLLM to process response $y$ as follows:
@@
Figure 3: Sparsification of s-MLLM. The VL Adaptor and Vision Encoder remain unchanged, while the LLM is upcycled from dense to sparse.
Sparsify s-MLLM. The principle of building our s-MLLM is downsizing the LLM while leaving the vision encoder and vision-language adaptor unmodified. To achieve this downsizing goal, we sparsify the dense s-MLLM by incorporating an MoE architecture. Specifically, Fig. 3 illustrates the process, where we apply the sparse upcycling technique (Komatsuzaki et al., 2022) to replicate $N$ feedforward networks (FFNs) as the expert modules. Additionally, we introduce a linear layer as the router, which dynamically activates the appropriate experts by predicting the probability of expert assignment. Given each token $x$ in the sequence, we first compute the routing value of $N$ experts:
@@ 3.2 Progressive Distillation
Our progressive distillation consists of two distinct stages, i.e., mimic distillation (Fig. 2 (a)) and preference distillation (Fig. 2 (b)). In the mimic distillation stage, the s-MLLM $\pi_{\text{S}}$ imitates the general and specific knowledge from the l-MLLM $\pi_{\text{T}}$. In the preference distillation stage, $\pi_{\text{S}}$ gains the preference knowledge of $\pi_{\text{T}}$ to further refine its output and reduce hallucinations. Both $\pi_{\text{S}}$ and $\pi_{\text{T}}$ are from the same LLM family. This ensures a consistent vocabulary space, which is essential for accurate imitation.
Initialization. Before distillation, we first align the vision encoder with the LLM via a learnable adapter, aiming to obtain a well-initialized dense version of $\pi_{\text{S}}$. The $\text{LLM}_{\phi}$ and $\text{ViT}_{\chi}$ are maintained frozen since their pre-trained parameters have already captured rich visual and language knowledge. Only $\text{Proj}_{\omega}$ is optimized to bridge the gap between the vision and language domain. For the initialization, we utilize common image-caption pairs from a widely used and curated dataset, which covers a diverse range of subjects and visual entities. The training objective is to minimize the cross-entropy of the generated tokens. The objective function is:
@@
Mimic Distillation. We decompose the comprehensive knowledge within $\pi_{\text{T}}$ into general and specialized aspects to address the challenges posed by their structural differences, which can complicate simultaneous learning. Subsequently, we conduct a general-to-specialized mimic distillation, which consists of two steps: dense-to-dense (D2D) and dense-to-sparse (D2S) distillation, to transfer the knowledge to $\pi_{\text{S}}$. This two-step approach balances the transfer of general and specialized knowledge through the progressive distillation, thereby enhancing overall performance. As illustrated in Fig. 3, we utilize the dense structure of $\pi_{\text{S}}$ during D2D to acquire general knowledge and transform it into a sparse structure during D2S to acquire complex specialized knowledge. Throughout both stages, $\pi_{\text{T}}$ remains unchanged.
a). Dense-to-Dense Distillation. In this stage, we aim to replicate the general knowledge of l-MLLM. Acquiring general knowledge first is crucial, as it establishes a broad foundation across various scenarios, enabling s-MLLM to develop a basic framework before advancing to multiple specialized tasks. To achieve this, we maintain $\text{ViT}_{\chi}$ frozen, and jointly optimize $\text{LLM}_{\phi}$ and $\text{Proj}_{\omega}$, with trainable parameters $\theta=\{\omega,\phi\}$. We leverage common image-caption pairs and conversation datasets. The training objective is to minimize the Kullback-Leibler divergence (KLD) between the output logits of s-MLLM and l-MLLM. The objective function is:
@@
where $V$ denotes the vocabulary, while $\pi_{\text{S}}(y_{k}\mid y_{<k},x)$ and $\pi_{\text{T}}(y_{k}\mid y_{<k},x)$ denote the probability of the predicted tokens for s-MLLM and l-MLLM.
b). Dense-to-Sparse Distillation. In this stage, our focus shifts to transfer the specialized knowledge of l-MLLM into s-MLLM to obtain advanced capabilities and achieve superior performance in complex tasks. However, directly learning this knowledge in the dense form of s-MLLM could be insufficient due to the diverse knowledge structure of different tasks. Therefore, we sparsify the dense s-MLLM by introducing multiple experts. As detailed in Section 3.1, we replicate $N$ feedforward networks (FFNs) within $\text{LLM}_{\phi}$ and add an MLP layer as a router, forming the experts with parameters $\phi_{e}$. This sparse architecture enables s-MLLM to activate the most relevant experts based on different inputs selectively, offering significant advantages in emulating the specialized knowledge of l-MLLM. For training, we leverage multi-task data, updating only the experts and the adaptor. We employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. Similar to the previous stage, we adopt the KLD as the training objective. Additionally, we include the standard next-token prediction objective, i.e., minimizing the cross-entropy loss of generated tokens by s-MLLM, to inject supervision from ground truth data which can reduce the existing bias in l-MLLM. The objective function is:
@@
Preference Distillation. In this stage, our goal is to distill the preference knowledge from l-MLLM to guide s-MLLM towards generating not only accurate but also reasonable responses, which is crucial in reducing hallucinations. For the training process, we effectively use preference data, which comprises meticulously paired positive responses $y^{+}$ and negative responses $y^{-}$ for the identical prompt $x$. Our preference distillation strategy is inspired by recent advancements in Preference Optimization (PO) (Rafailov et al., 2024; Ethayarajh et al., 2024), which bypasses the need for training a reward model by directly training on an offline preference dataset. Our key insight is to treat l-MLLM as the reference model to provide insights on what constitutes “good” and “bad”, thereby establishing a fundamental reference for s-MLLM.
Specifically, the training objective is to optimize s-MLLM to assign higher probabilities to positive responses and lower probabilities to negative ones compared to l-MLLM. This training process involves two key optimization aspects: First, s-MLLM aims to align with the teacher model in distinguishing positive from negative responses. Second, s-MLLM seeks to surpass l-MLLM by assigning higher probabilities to positive responses and lower probabilities to negative responses. We only train the experts and adaptor in s-MLLM and employ a Top-$k$ routing strategy to select the experts. The trainable parameters are $\theta=\{\omega,\phi_{e}\}$. The objective function is:
@@
where $\pi_{\text{S}}(y^{+}\mid x)$ and $\pi_{\text{S}}(y^{-}\mid x)$ denote the probabilities of positive and negative responses for s-MLLM, and $\pi_{\text{T}}(y^{+}\mid x)$ and $\pi_{\text{T}}(y^{-}\mid x)$ denote the same for l-MLLM.